# Locally rigid $\infty$-categories

### v1

@@ 3 Trace-class maps
Recall from Example 2.9 that for $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}})$, trace-class maps are the same thing as $\mathcal{V}$-atomic maps in $\mathcal{V}$. One of the key features of trace-class maps is the following:
@@
This is formally analogous to the diagonal filler appearing in the definition of compact maps, however if we think of tensoring as analogous to colimits, here the assumption is on “a type of colimit diagram”, while in the definition of compact maps, the assumption is on “$m$”. It is not clear to the author whether there is a common framework for these two statements; however they have the same kind of underlying “meaning” (one can commute things that ought not to commute), and the same kind of use.
@@ 4 Local rigidity and rigidification
How precisely can we compare ${\mathrm{at}}_{\mathcal{M}}^{\mathcal{V}}(x,y)$ and ${\mathrm{at}}_{\mathcal{M}}^{\mathcal{W}}(x,y)$
@@
It turns out that these two conditions (being smooth and proper) are implied by the stronger condition of being rigid. This notion will also be of help to study our other questions; as well as its variant, that of a locally rigid $\mathcal{V}$-algebra. We point out that these notions have been introduced by Gaitsory and Rozenblyum and further studied in [GR19b, HSS17] and [AGK+20].
@@ 4.1 Local rigidity
Assume now that every $\mathcal{V}$-atomic object in $\mathcal{W}$ is dualizable. In this case, the multiplication map sends pairs of $\mathcal{V}$-atomics to $\mathcal{V}$-atomics (because a $\mathcal{V}$-atomic tensored with a $\mathcal{W}$-dualizable is $\mathcal{V}$-atomic, by an argument similar to [Ram24, Example 1.23]), and thus is an internal left adjoint in $\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}})$. We are left with proving that its right adjoint $m^{R}:\mathcal{W}\to\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}$ is strictly $\mathcal{W}$-bilinear. Since it is $\mathcal{V}$-linear and colimit-preserving, and since $\mathcal{W}$ is $\mathcal{V}$-atomically generated, it suffices to check that the projection formula $w\otimes m^{R}(w^{\prime})\to m^{R}(w\otimes w^{\prime})$ is an equivalence for $\mathcal{V}$-atomic $w$. But now by assumption, $w$ is dualizable and so this map is an equivalence by 3.8.
@@
For the converse, note that $-\otimes_{\mathcal{W}}\mathcal{M}:\mathbf{Mod}_{\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}}(\mathbf{Pr}^{\mathrm{L}})\to\mathbf{Mod}_{\mathcal{W}}(\mathbf{Pr}^{\mathrm{L}})$ is a $2$-functor, so it preserves internal left adjoints, which allows us to conclude. ∎
@@
where we have used 2.13.
@@
Finally, using 2.21 (since the functor $\mathbf{1}_{\mathcal{W}}\otimes\hat{j}$ is a fully faithful $\mathcal{W}$-internal left adjoint), the latter is equivalent to ${\mathrm{at}}^{\mathcal{W}}_{\mathcal{W}\otimes_{\mathcal{W}}\mathcal{M}}(\mathbf{1}_{\mathcal{W}}\otimes x,\mathbf{1}_{\mathcal{W}}\otimes y)$, and it is easy to check that all of this is happening over the relevant $\hom$-objects, so that this proves the claim. ∎
@@
Combine the previous lemma with 4.11 and Example 2.20. ∎
@@
So first assume that $\mathcal{W}$ is locally rigid. As it is dualizable, any dualizable $\mathcal{W}$-module is $\mathcal{V}$-dualizable [Lur12, Proposition 4.6.4.4.]. Now if $\mathcal{M}$ is dualizable over $\mathcal{W}$ and $\alpha$ is $\mathbf{Sp}$-atomic in $\mathcal{M}$, then by 4.13, $\mathbf{1}_{\mathcal{W}}\boxtimes\alpha$ is $\mathcal{W}$-atomic in $\mathcal{W}\otimes\mathcal{M}$. By 4.11, the counit $\mathcal{W}\otimes\mathcal{M}\to\mathcal{M}$ is an internal left adjoint and so it sends $\mathcal{W}$-atomic maps to $\mathcal{W}$-atomic maps, so that $\alpha$ is $\mathcal{W}$-atomic too.
Conversely, suppose $\mathcal{W}$ is $\mathbf{Sp}$-dualizable, and $\mathbf{Sp}$-atomic maps in $\mathcal{W}$ are trace-class. We have to prove that to $\mathcal{W}\otimes\mathcal{W}\to\mathcal{W}$ is a $\mathcal{W}$-internal left adjoint. First note that it sends pairs $f\boxtimes g$ of $\mathbf{Sp}$-atomic maps to the tensor product of a $\mathbf{Sp}$-atomic map and a trace-class map (by assumption) which is thus $\mathbf{Sp}$-atomic by 3.9. Thus it sends enough $\mathbf{Sp}$-atomic maps to $\mathbf{Sp}$-atomic maps, and is thus a $\mathbf{Sp}$-internal left adjoint by [Ram24, Addendum 2.42].
We now prove that the right adjoint $\mu^{R}$ is actually $\mathcal{W}$-linear on (say) the left. For this, we use 3.7. Writing any object of $\mathcal{W}$ as a colimit of compactly exhaustible objects, we see that $f^{R}$ is $\mathcal{W}$-linear. ∎
@@
Thus the colimit-closure of $\mathcal{W}$-atomic telescopes in $\mathcal{M}$ is the whole of $\mathcal{M}$, and thus by 2.25 $\mathcal{M}$ is dualizable over $\mathcal{W}$. ∎
@@
Note that $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{M}$ is dualizable over $\mathcal{W}$, and by 4.11, the action map $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{M}\to\mathcal{M}$ is an internal left adjoint. Its right adjoint is clearly conservative, so we may conclude by [Ram24, Proposition 1.41]. ∎
@@
We have maps $\mathbf{1}_{\mathcal{W}}\otimes{\mathrm{at}}_{\mathcal{M}}^{\mathcal{V}}(-,y)\to{\mathrm{at}}_{\mathcal{M}}^{\mathcal{W}}(-,y)$ for all $y\in\mathcal{M}$ by 4.14. It follows from this together with 1.12 that $\mathcal{V}$-atomic presentations in $\mathcal{M}$ are also $\mathcal{W}$-atomic presentations, and so by 2.25, $\mathcal{M}$ is dualizable over $\mathcal{W}$. ∎
@@
This is the same argument as the end of the proof of 4.15. ∎
@@
The same argument as in 4.17 works, as the action map $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{M}\to\mathcal{M}$ is natural in the input, and we can use [Ram24, Lemma 1.30] in place of [Ram24, Proposition 1.41]. ∎
@@
This is essentially the same argument as in the case of $\mathbf{Sp}$. More precisely, using 4.14 and 1.12, we may use 2.31 with the weights $W_{s}$ being of the form $\mathbf{1}_{\mathcal{W}}\otimes{\mathrm{at}}_{\mathcal{M}}^{\mathcal{V}}(-,x_{s})$, together with the fact that these are preserved by Example 2.20. ∎
@@
Fix $\mathcal{M}$ a $\mathcal{W}$-module, by 4.11 it suffices to prove that the action map $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{M}\to\mathcal{M}$ is a $\mathcal{W}$-internal left adjoint. $\mathcal{W}$ is generated under colimits by the images of the $\mathcal{W}_{i}$’s, so to check that it is an internal left adjoint, it suffices to check that it is a $\mathcal{V}$-internal left adjoint, and a $\mathcal{W}_{i}$-internal left adjoint for all $i$.
For the former, we note that it is a colimit along internal left adjoints of $\mathcal{W}_{i}\otimes_{\mathcal{V}}\otimes\mathcal{M}\to\mathcal{M}$, each of which is an internal left adjoint as $\mathcal{W}_{i}$ is locally rigid.
The latter works the same, noting now that for each $i$, the forgetful map $I_{i/}\to I$ is cofinal by siftedness of $I$, so that we can instead assume we had a diagram of locally rigid $\mathcal{W}_{i}$-algebras along $\mathcal{W}_{i}$-internal left adjoints by 4.18. ∎
@@
We will study further categorical properties of the category of locally rigid categories in Section 5.2.
We now give an a posteriori description of the duality datum of a locally rigid category. Recall that dualizability is assumed in the definition of local rigidity [^9] In contrast to the case of rigidity where it could be omitted, cf. 4.54.. However, once we have assumed it, the duality datum has a simple description :
@@
[…]
@@
One can use the rigid case (see 4.54) as well as the theory of rigidification (see 4.80) to give an alternative proof. We will describe this alternative proof later in Section 4.3. We will also see there that $\Gamma_{!}$ is quite explicit: it is given by ${\mathrm{at}}_{\mathcal{W}}^{\mathcal{V}}(\mathbf{1}_{\mathcal{W}},-)$; see 4.63, 4.64 and 4.65. ∎
@@
Question 4.25 (Local compactness).
@@
[…]
@@
We claim that there is a dashed arrow filling this diagram, namely the composite $\hom_{\mathcal{M}}(x,y)=\hom_{\mathcal{M}}(x,y)\otimes\mathbf{1}_{\mathcal{W}}\xrightarrow{\mathrm{id}\otimes\hat{g}}\hom_{\mathcal{M}}(x,y)\otimes{\mathrm{at}}_{\mathcal{M}}(y,z)\to{\mathrm{at}}_{\mathcal{M}}(x,z)$, where the last map is from 2.17. That the inner triangles commute also follows from loc. cit..
@@
It is unclear to the author how optimal this is. It is very likely that with some good enough understanding of the posets $\omega_{n}$, as $n$ varies, one can prove with similar techniques a similar lemma involving composites of $n+1$ locally compact maps. But it is also possible that all of this is inessential, and that the answer to Question 4.25 is simply “yes”, or maybe yes with composites of two locally compact maps with no set-theoretic assumptions.
@@ 4.2 Rigidity
This follows from the definition and from Example 4.6. ∎
@@
We have seen in Example 4.9 that it is locally rigid. Since $L:\mathcal{W}\to\mathcal{W}_{L}$ is a $\mathcal{V}$-internal left adjoint, it sends the $\mathcal{V}$-atomic object $\mathbf{1}_{\mathcal{W}}$ to a $\mathcal{V}$-atomic object, namely $\mathbf{1}_{\mathcal{W}_{L}}$. ∎
@@
By Example 4.12, it is locally rigid. Furthermore, $\mathcal{V}\to\mathcal{W}$ is a $\mathcal{V}$-internal left adjoint as $\mathbf{1}_{\mathcal{W}}$ is $\mathcal{V}$-atomic; $\mathcal{W}\to\mathcal{U}$ is a $\mathcal{W}$-internal left adjoint as $\mathbf{1}_{\mathcal{U}}$ is $\mathcal{W}$-atomic, but now this immediately implies that it is als a $\mathcal{V}$-internal left adjoint, so the composite $\mathcal{V}\to\mathcal{U}$ also is, thus proving that $\mathbf{1}_{\mathcal{U}}$ is $\mathcal{V}$-atomic.
@@
More generally, assume $\mathcal{W}_{0}\to\mathcal{W}_{1}$ is a symmetric-monoidal colimit-preserving functor between commutative $\mathcal{V}$-algebras, where $\mathcal{W}_{0}$ is locally rigid and the unit in $\mathcal{W}_{1}$ is $\mathcal{V}$-atomic. The latter condition implies, by tensoring with $\mathcal{W}_{0}$, that $\mathcal{W}_{0}\to\mathcal{W}_{0}\otimes_{\mathcal{V}}\mathcal{W}_{1}$ is a $\mathcal{W}_{0}$-internal left adjoint; and 4.11 implies that $\mathcal{W}_{0}\otimes_{\mathcal{V}}\mathcal{W}_{1}\to\mathcal{W}_{1}$ is also a $\mathcal{W}_{0}$-internal left adjoint. It follows that the map $\mathcal{W}_{0}\to\mathcal{W}_{1}$ is a $\mathcal{W}_{0}$-internal left adjoint, and thus by restricting scalars, a $\mathcal{V}$-internal left adjoint.
@@
We begin with the locally rigid situation.
First, assume $\mathcal{W}$ is locally rigid. In this case, the identity of a $\mathcal{V}$-atomic object in $\mathcal{W}$ is a $\mathcal{V}$-atomic map, hence by local rigidity a $\mathcal{W}$-atomic map, i.e. a trace-class map, and hence the object is dualizable.
Conversely, assume all $\mathcal{V}$-atomics are dualizable. In that case, $\mathcal{W}$ is generated under colimits and $\mathcal{V}$-tensors by dualizables and so any $\mathcal{W}$-linear functor, $\mathcal{V}$-internally left adjoint is a $\mathcal{W}$-internal left adjoint. It thus suffices to prove that the multiplication map $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}\to\mathcal{W}$ is a $\mathcal{V}$-internal left adjoint. But its source is atomically generated, and it sends enough $\mathcal{V}$-atomics to $\mathcal{V}$-atomics (because a $\mathcal{V}$-atomic tensor a dualizable is $\mathcal{V}$-atomic).
Now the rigid situation. Clearly, if $\mathcal{W}$ is rigid, by the above, all $\mathcal{V}$-atomics are dualizable. The converse of course holds if $\mathbf{1}_{\mathcal{W}}$ is atomic.
Conversely, assume atomics and dualizables agree. It follows from the above that $\mathcal{W}$ is locally rigid, and because $\mathbf{1}_{\mathcal{W}}$ is atomic, it follows that $\mathcal{W}$ is rigid. ∎
@@
In 4.42 and 4.14, we have constructed maps in both directions lying over $\hom_{\mathcal{M}}^{\mathcal{V}}(x,-)$. Note that this is already sufficient to prove that the notions of $\mathcal{V}$-atomic and $\mathcal{W}$-atomic maps coincide. Now we prove the more precise statement. Fix $\kappa$ large enough.
@@
The functor ${\mathrm{at}}_{\mathcal{M}}^{\mathcal{W}}(x,-)$ is $\mathcal{W}$-linear, so by 4.55 the top map is of the form $f^{R}(\alpha)$ for some $\alpha:{\mathrm{at}}^{\mathcal{W}}_{\mathcal{M}}(x,-)\to{\mathrm{at}}^{\mathcal{W}}_{\mathcal{M}}(x,-)$. The point now is to extend the conclusion 4.55 to $\mathcal{V}$-enriched functors.
@@
where the second equivalence comes from 1.1 and the third from 4.55.
@@
From this perspective, unwinding the previous equivalences, we find that the map given by postcomposing with $f^{R}$ is identified with $j^{*}:\operatorname{Fun}^{L}_{\mathcal{V}}(\mathcal{P}_{\mathcal{W}}(\mathcal{M}^{\kappa}),\mathcal{V})\to\operatorname{Fun}^{L}_{\mathcal{V}}(\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa}),\mathcal{V})$, where $j:\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})\to\mathcal{P}_{\mathcal{W}}(\mathcal{M}^{\kappa})$ is the $\mathcal{V}$-linear map induced by the embedding $\mathcal{V}$-enriched $\mathcal{M}^{\kappa}\to\mathcal{P}_{\mathcal{W}}(\mathcal{M}^{\kappa})$, using 1.1.
@@
By 4.48, any map between rigid $\mathcal{V}$-algebras is an internal left adjoint. Thus 4.19 already implies that any such colimit is locally rigid.

### v2

@@ 3 Trace-class maps
Recall from Section˜2 that for $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}})$, trace-class maps are the same thing as $\mathcal{V}$-atomic maps in $\mathcal{V}$. One of the key features of trace-class maps is the following:
@@
This is formally analogous to the diagonal filler appearing in the definition of compact maps, however if we think of tensoring as analogous to colimits, here the assumption is on “a type of colimit diagram”, while in the definition of compact maps, the assumption is on “mm”. It is not clear to the author whether there is a common framework for these two statements; however they have the same kind of underlying “meaning” (one can commute things that ought not to commute), and the same kind of use.
@@ 4 Local rigidity and rigidification
How precisely can we compare ${\mathrm{at}}_{\mathcal{M}}^{\mathcal{V}}(x,y)$ and ${\mathrm{at}}_{\mathcal{M}}^{\mathcal{W}}(x,y)$ ?
@@
It turns out that these two conditions (being smooth and proper) are implied by the stronger condition of being rigid. This notion will also be of help to study our other questions; as well as its variant, that of a locally rigid $\mathcal{V}$-algebra. We point out that these notions have been introduced by Gaitsory and Rozenblyum and further studied in [GR19a, HSS17] and [AGK+20].
@@ 4.1 Local rigidity
Assume now that every $\mathcal{V}$-atomic object in $\mathcal{W}$ is dualizable. In this case, the multiplication map sends pairs of $\mathcal{V}$-atomics to $\mathcal{V}$-atomics (because a $\mathcal{V}$-atomic tensored with a $\mathcal{W}$-dualizable is $\mathcal{V}$-atomic, by an argument similar to [RAM24, Example 1.23]), and thus is an internal left adjoint in $\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}})$. We are left with proving that its right adjoint $m^{R}:\mathcal{W}\to\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}$ is strictly $\mathcal{W}$-bilinear. Since it is $\mathcal{V}$-linear and colimit-preserving, and since $\mathcal{W}$ is $\mathcal{V}$-atomically generated, it suffices to check that the projection formula $w\otimes m^{R}(w^{\prime})\to m^{R}(w\otimes w^{\prime})$ is an equivalence for $\mathcal{V}$-atomic $w$. But now by assumption, $w$ is dualizable and so this map is an equivalence by Section˜3.
@@
For the converse, note that $\otimes_{\mathcal{W}}\mathcal{M}:\mathbf{Mod}_{\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}}(\mathbf{Pr}^{\mathrm{L}})\to\mathbf{Mod}_{\mathcal{W}}(\mathbf{Pr}^{\mathrm{L}})$ is a $2$-functor, so it preserves internal left adjoints, which allows us to conclude. ∎
@@
where we have used Section˜2.
@@
Finally, using Section˜2 (since the functor $\mathbf{1}_{\mathcal{W}}\otimes\hat{j}$ is a fully faithful $\mathcal{W}$-internal left adjoint), the latter is equivalent to ${\mathrm{at}}^{\mathcal{W}}_{\mathcal{W}\otimes_{\mathcal{W}}\mathcal{M}}(\mathbf{1}_{\mathcal{W}}\otimes x,\mathbf{1}_{\mathcal{W}}\otimes y)$, and it is easy to check that all of this is happening over the relevant $\hom$-objects, so that this proves the claim. ∎
@@
Combine the previous lemma with Section˜4.1 and Section˜2. ∎
@@
So first assume that $\mathcal{W}$ is locally rigid. As it is dualizable, any dualizable $\mathcal{W}$-module is $\mathcal{V}$-dualizable [LUR12, Proposition 4.6.4.4.]. Now if $\mathcal{M}$ is dualizable over $\mathcal{W}$ and $\alpha$ is $\mathbf{Sp}$-atomic in $\mathcal{M}$, then by Section˜4.1, $\mathbf{1}_{\mathcal{W}}\boxtimes\alpha$ is $\mathcal{W}$-atomic in $\mathcal{W}\otimes\mathcal{M}$. By Section˜4.1, the counit $\mathcal{W}\otimes\mathcal{M}\to\mathcal{M}$ is an internal left adjoint and so it sends $\mathcal{W}$-atomic maps to $\mathcal{W}$-atomic maps, so that $\alpha$ is $\mathcal{W}$-atomic too.
Conversely, suppose $\mathcal{W}$ is $\mathbf{Sp}$-dualizable, and $\mathbf{Sp}$-atomic maps in $\mathcal{W}$ are trace-class. We have to prove that to $\mathcal{W}\otimes\mathcal{W}\to\mathcal{W}$ is a $\mathcal{W}$-internal left adjoint. First note that it sends pairs $f\boxtimes g$ of $\mathbf{Sp}$-atomic maps to the tensor product of a $\mathbf{Sp}$-atomic map and a trace-class map (by assumption) which is thus $\mathbf{Sp}$-atomic by Section˜3. Thus it sends enough $\mathbf{Sp}$-atomic maps to $\mathbf{Sp}$-atomic maps, and is thus a $\mathbf{Sp}$-internal left adjoint by [RAM24, Addendum 2.42].
We now prove that the right adjoint $\mu^{R}$ is actually $\mathcal{W}$-linear on (say) the left. For this, we use Section˜3. Writing any object of $\mathcal{W}$ as a colimit of compactly exhaustible objects, we see that $f^{R}$ is $\mathcal{W}$-linear. ∎
@@
Thus the colimit-closure of $\mathcal{W}$-atomic telescopes in $\mathcal{M}$ is the whole of $\mathcal{M}$, and thus by ˜2.25 $\mathcal{M}$ is dualizable over $\mathcal{W}$. ∎
@@
Note that $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{M}$ is dualizable over $\mathcal{W}$, and by Section˜4.1, the action map $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{M}\to\mathcal{M}$ is an internal left adjoint. Its right adjoint is clearly conservative, so we may conclude by [RAM24, Proposition 1.41]. ∎
@@
We have maps $\mathbf{1}_{\mathcal{W}}\otimes{\mathrm{at}}_{\mathcal{M}}^{\mathcal{V}}(-,y)\to{\mathrm{at}}_{\mathcal{M}}^{\mathcal{W}}(-,y)$ for all $y\in\mathcal{M}$ by Section˜4.1. It follows from this together with Section˜1 that $\mathcal{V}$-atomic presentations in $\mathcal{M}$ are also $\mathcal{W}$-atomic presentations, and so by ˜2.25, $\mathcal{M}$ is dualizable over $\mathcal{W}$. ∎
@@
This is the same argument as the end of the proof of Section˜4.1. ∎
@@
The same argument as in Section˜4.1 works, as the action map $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{M}\to\mathcal{M}$ is natural in the input, and we can use [RAM24, Lemma 1.30] in place of [RAM24, Proposition 1.41]. ∎
@@
This is essentially the same argument as in the case of $\mathbf{Sp}$. More precisely, using Section˜4.1 and Section˜1, we may use Section˜2 with the weights $W_{s}$ being of the form $\mathbf{1}_{\mathcal{W}}\otimes{\mathrm{at}}_{\mathcal{M}}^{\mathcal{V}}(-,x_{s})$, together with the fact that these are preserved by Section˜2. ∎
We can use these results to prove a converse to Section˜4.1:
Corollary 4.19.
Let $f:\mathcal{V}\to\mathcal{W},g:\mathcal{W}\to\mathcal{U}$ two morphisms in $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}})$. If $f$ and $gf$ are locally rigid, then $g$ is also locally rigid.
Remark 4.20.
We do not have a full 22-out-of-33 for local rigidity, as the example $\mathbf{Sp}\to\operatorname{Fun}(\mathbb{N},\mathbf{Sp})\to\mathbf{Sp}$ shows.
Proof.
By Section˜4.1, $\mathcal{U}$ is dualizable over $\mathcal{W}$. Consider now the multiplication map $\mathcal{U}\otimes_{\mathcal{W}}\mathcal{U}\to\mathcal{U}$, which we need to show is a $\mathcal{U}$-internal left adjoint on either side. Note that the map $\mathcal{U}\otimes_{\mathcal{V}}\mathcal{U}\to\mathcal{U}\otimes_{\mathcal{W}}\mathcal{U}$ is a map of the form $\mathcal{M}\to\mathcal{M}\otimes_{\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}}\mathcal{W}$ and so by local rigidity of $\mathcal{W}$, it is a $\mathcal{W}$-internal left adjoint, and hence a $\mathcal{V}$-internal left adjoint. Furthermore, its image generates the target under colimits. Thus, by [RAM24, Lemma 1.30] and local rigidity of $gf$, the map $\mathcal{U}\otimes_{\mathcal{W}}\mathcal{U}\to\mathcal{U}$ is a $\mathcal{V}$-internal left adjoint. By Section˜4.1, it is a $\mathcal{U}$-internal left adjoint, as was to be shown. ∎
@@
Fix $\mathcal{M}$ a $\mathcal{W}$-module, by Section˜4.1 it suffices to prove that the action map $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{M}\to\mathcal{M}$ is a $\mathcal{W}$-internal left adjoint. $\mathcal{W}$ is generated under colimits by the images of the $\mathcal{W}_{i}$’s, so to check that it is an internal left adjoint, it suffices to check that it is a $\mathcal{V}$-internal left adjoint, and a $\mathcal{W}_{i}$-internal left adjoint for all $i$.
For the former, we note that it is a colimit along internal left adjoints of $\mathcal{W}_{i}\otimes_{\mathcal{V}}\mathcal{M}\to\mathcal{M}$, each of which is an internal left adjoint as $\mathcal{W}_{i}$ is locally rigid.
The latter works the same, noting now that for each $i$, the forgetful map $I_{i/}\to I$ is cofinal by siftedness of $I$, so that we can instead assume we had a diagram of locally rigid $\mathcal{W}_{i}$-algebras along $\mathcal{W}_{i}$-internal left adjoints by Section˜4.1. ∎
@@
We will study further categorical properties of the category of locally rigid categories in Section˜5.2.
We now give an a posteriori description of the duality datum of a locally rigid category. Recall that dualizability is assumed in the definition of local rigidity [^10] In contrast to the case of rigidity where it could be omitted, cf. Section 4.2.. However, once we have assumed it, the duality datum has a simple description :
@@
$\mathcal{V}\xrightarrow{\eta}\mathcal{W}\xrightarrow{\mu^{R}}\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W},\qquad\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}\xrightarrow{\mu}\mathcal{W}\xrightarrow{\Gamma_{!}}\mathcal{V}$
@@
One can use the rigid case (see Section˜4.2) as well as the theory of rigidification (see Section˜4.3.1) to give an alternative proof. We will describe this alternative proof later in Section˜4.3. We will also see there that $\Gamma_{!}$ is quite explicit: it is given by ${\mathrm{at}}_{\mathcal{W}}^{\mathcal{V}}(\mathbf{1}_{\mathcal{W}},-)$; see ˜4.67, Section˜4.3 and Section˜4.3. ∎
@@
Question 4.27 (Local compactness).
@@
Let $x\xrightarrow{f}y\xrightarrow{g}z$ be two locally compact maps in $\mathcal{M}$.
@@
We claim that there is a dashed arrow filling this diagram, namely the composite $\hom_{\mathcal{M}}(x,y)=\hom_{\mathcal{M}}(x,y)\otimes\mathbf{1}_{\mathcal{W}}\xrightarrow{\mathrm{id}\otimes\hat{g}}\hom_{\mathcal{M}}(x,y)\otimes{\mathrm{at}}_{\mathcal{M}}(y,z)\to{\mathrm{at}}_{\mathcal{M}}(x,z)$, where the last map is from Section˜2. That the inner triangles commute also follows from loc. cit..
@@
It is unclear to the author how optimal this is. It is very likely that with some good enough understanding of the posets $\omega_{n}$, as $n$ varies, one can prove with similar techniques a similar lemma involving composites of $n+1$ locally compact maps. But it is also possible that all of this is inessential, and that the answer to Section˜4.1 is simply “yes”, or maybe yes with composites of two locally compact maps with no set-theoretic assumptions.
@@ 4.2 Rigidity
This follows from the definition and from Section˜4.1. ∎
@@
We have seen in Section˜4.1 that it is locally rigid. Since $L:\mathcal{W}\to\mathcal{W}_{L}$ is a $\mathcal{V}$-internal left adjoint, it sends the $\mathcal{V}$-atomic object $\mathbf{1}_{\mathcal{W}}$ to a $\mathcal{V}$-atomic object, namely $\mathbf{1}_{\mathcal{W}_{L}}$. ∎
@@
By Section˜4.1, it is locally rigid. Furthermore, $\mathcal{V}\to\mathcal{W}$ is a $\mathcal{V}$-internal left adjoint as $\mathbf{1}_{\mathcal{W}}$ is $\mathcal{V}$-atomic; $\mathcal{W}\to\mathcal{U}$ is a $\mathcal{W}$-internal left adjoint as $\mathbf{1}_{\mathcal{U}}$ is $\mathcal{W}$-atomic, but now this immediately implies that it is als a $\mathcal{V}$-internal left adjoint, so the composite $\mathcal{V}\to\mathcal{U}$ also is, thus proving that $\mathbf{1}_{\mathcal{U}}$ is $\mathcal{V}$-atomic.
@@
Here, we can directly prove a weak converse:
Corollary 4.46.
Let $f:\mathcal{V}\to\mathcal{W},g:\mathcal{W}\to\mathcal{U}$ be two morphisms in $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}})$. If $f$ and $gf$ are rigid, so is $g$.
Proof.
By Section˜4.1, $g$ is locally rigid, so it suffices to prove that $\mathbf{1}_{\mathcal{U}}$ is $\mathcal{W}$-atomic. Since it is $\mathcal{V}$-atomic, this follows from Section˜4.1. ∎
@@
More generally, assume $\mathcal{W}_{0}\to\mathcal{W}_{1}$ is a symmetric-monoidal colimit-preserving functor between commutative $\mathcal{V}$-algebras, where $\mathcal{W}_{0}$ is locally rigid and the unit in $\mathcal{W}_{1}$ is $\mathcal{V}$-atomic. The latter condition implies, by tensoring with $\mathcal{W}_{0}$, that $\mathcal{W}_{0}\to\mathcal{W}_{0}\otimes_{\mathcal{V}}\mathcal{W}_{1}$ is a $\mathcal{W}_{0}$-internal left adjoint; and Section˜4.1 implies that $\mathcal{W}_{0}\otimes_{\mathcal{V}}\mathcal{W}_{1}\to\mathcal{W}_{1}$ is also a $\mathcal{W}_{0}$-internal left adjoint. It follows that the map $\mathcal{W}_{0}\to\mathcal{W}_{1}$ is a $\mathcal{W}_{0}$-internal left adjoint, and thus by restricting scalars, a $\mathcal{V}$-internal left adjoint.
@@
The locally rigid case is Section˜4.2.
Now if $\mathcal{W}$ is rigid then by the above, all $\mathcal{V}$-atomics are dualizable. The converse of course holds if $\mathbf{1}_{\mathcal{W}}$ is atomic.
Conversely, assume atomics and dualizables agree. It follows again from the above that $\mathcal{W}$ is locally rigid, and because $\mathbf{1}_{\mathcal{W}}$ is atomic, it follows that $\mathcal{W}$ is rigid. ∎
@@
In Section˜4.2 and Section˜4.1, we have constructed maps in both directions lying over $\hom_{\mathcal{M}}^{\mathcal{V}}(x,-)$. Note that this is already sufficient to prove that the notions of $\mathcal{V}$-atomic and $\mathcal{W}$-atomic maps coincide. Now we prove the more precise statement. Fix $\kappa$ large enough.
@@
The functor ${\mathrm{at}}_{\mathcal{M}}^{\mathcal{W}}(x,-)$ is $\mathcal{W}$-linear, so by Section˜4.2 the top map is of the form $f^{R}(\alpha)$ for some $\alpha:{\mathrm{at}}^{\mathcal{W}}_{\mathcal{M}}(x,-)\to{\mathrm{at}}^{\mathcal{W}}_{\mathcal{M}}(x,-)$. The point now is to extend the conclusion Section˜4.2 to $\mathcal{V}$-enriched functors.
@@
where the second equivalence comes from ˜1.1 and the third from Section˜4.2.
@@
From this perspective, unwinding the previous equivalences, we find that the map given by postcomposing with $f^{R}$ is identified with $j^{*}:\operatorname{Fun}^{L}_{\mathcal{V}}(\mathcal{P}_{\mathcal{W}}(\mathcal{M}^{\kappa}),\mathcal{V})\to\operatorname{Fun}^{L}_{\mathcal{V}}(\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa}),\mathcal{V})$, where $j:\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})\to\mathcal{P}_{\mathcal{W}}(\mathcal{M}^{\kappa})$ is the $\mathcal{V}$-linear map induced by the embedding $\mathcal{V}$-enriched $\mathcal{M}^{\kappa}\to\mathcal{P}_{\mathcal{W}}(\mathcal{M}^{\kappa})$, using ˜1.1.
@@
By Section˜4.2, any map between rigid $\mathcal{V}$-algebras is an internal left adjoint. Thus Section˜4.1 already implies that any such colimit is locally rigid.